OpenStudy (anonymous):

Write a rule for the linear function whose graph has slope m and contains the given point. m = 3; (-2, -2) a. y = 3x + 4 c. y = 3x + 2 b. y = 3x - 4 d. y = 3x - 2

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

y=3x+b... b=y-3x... let x=-2 let y=-2 and (x=y) b=(-2)-3(-2) b=(-2+6) b=6-2 b=4. hence, y=mx+b becomes... m=-3 b=4 becomes... Ans. y=-3x+4.....! y=3x-4! which also equals y=3x-4.... hence the answer is...